OpenStudy (anonymous):

Evaluate b2c-1 for b = -4 and c = 2.

OpenStudy (anonymous):

-17

OpenStudy (anonymous):

-4 x 2 x 2 - 1

OpenStudy (anonymous):

just plug in the values.

jigglypuff314 (jigglypuff314):

I believe you meant for the numbers to be exponents if so, (-4)^2 * (2)^-1

OpenStudy (anonymous):

so b^2(c)-1 =(-4)^2(2)-1 =16(2)-1 =32-1 =31

OpenStudy (anonymous):

If it was exponents. If not, it'd be -17

OpenStudy (anonymous):

yeah. they weren't clear on that

jigglypuff314 (jigglypuff314):

usually when the numbers are spaced like that, it would suggest exponents (-4)^2 * (2)^-1 would not be 31 though

OpenStudy (anonymous):

These are the answers, -8 4 8 32 and sorry the numbers are supposed to be exponents.

OpenStudy (anonymous):

is "c" in the exponents?

OpenStudy (anonymous):

I think its 32 though.

jigglypuff314 (jigglypuff314):

alright (-4)^2 * (2)^-1 expanded becomes (-4)*(-4)*(1/2) = ? can you do that

OpenStudy (anonymous):

no, c=2

OpenStudy (anonymous):

@jigglypuff314 8

jigglypuff314 (jigglypuff314):

correct :)

OpenStudy (anonymous):

do you mean that c is to the power of -1? or just minus 1?

OpenStudy (anonymous):

So it equals 8?

jigglypuff314 (jigglypuff314):

if c is to the power of -1 then yes, 8

OpenStudy (anonymous):



OpenStudy (anonymous):

is how it looks dg, thank you jigglypuff